The perimeter of a rectangle is twie as long as it is wide is 90 meters. finds the area?
Answer by nerdybill(7384)

The perimeter of a rectangle is twie as long as it is wide is 90 meters. finds the area?
begin by finding the width and length:
Let w = width
then
2w = length
.
2(w + 2w) = 90
2(3w) = 90
3w = 45
w = 15 meters
.
length:
2w = 2(15) = 30 meters
.
Area then is:
15*30 = 450 square meters